If α+β=π/3 the and sin α=1/2 then find the value of sinβ
what is the answer for above question

Answer by stanbon(75887)

If α+β=π/3 the and sin α=1/2 then find the value of sinβ
what is the answer for above question
------
If sin(a) = 1/2, a = pi/6
------
Equation:
a + b = pi/3
(pi/6) + b = pi/3
-----
pi/6 + b = (2pi/6)
------
b = pi/6
------------
Ans: sin(b) = 1/2
=======================
Cheers,
Stan H.